OpenStudy (mathmath333):

A person draws a card from a pack of 52, replaces it and shuffles it. He continues doing so until he draws a heart. What is the probablity that he has to make at least 3 trials?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{A person draws a card from a pack of 52, replaces it} \hspace{.33em}\\~\\ & \normalsize \text{ and shuffles it. He continues doing so until he draws }\hspace{.33em}\\~\\ & \normalsize \text{a heart. What is the probablity that he has to make at}\hspace{.33em}\\~\\ & \normalsize \text{least 3 trials?}\hspace{.33em}\\~\\ & a). \dfrac{3}{17} \hspace{.33em}\\~\\ & b). \dfrac{8}{19} \hspace{.33em}\\~\\ & c). \dfrac{2}{17} \hspace{.33em}\\~\\ & d). \dfrac{11}{17} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (phi):

this sounds like the "hit or miss" problem from yesterday

OpenStudy (mathmath333):

u mean binomuial distrbtion

OpenStudy (mathmath333):

hint is given: the player didnt get a heart in first 2 cases

OpenStudy (mathmath333):

i think by the hint it can be 1-(3/4)^2=7/16 but that is not in options

OpenStudy (mathmath333):

answer given is d.)11/16

OpenStudy (phi):

doing a tree